# Lattice paths and the Geode

### v1

@@ Abstract.
Let $t_{1},t_{2},\dots$ be variables, and let $S$ be the formal power series in the variables $t_{1},t_{2},\dots$ satisfying $S=1+\sum_{i=1}^{\infty}t_{n}S^{n}.$ Let $S_{1}=\sum_{n=1}^{\infty}t_{n}$. Wildberger and Rubine recently showed that there is a formal power series $G$ in the $t_{i}$, which they called the “Geode,” satisfying $S=1+GS_{1}$. In this paper we discuss some of the properties of the Geode and of the related series $H=G/S$, which satisfies $S=1/(1-HS_{1})$. We show that
@@ 1. Introduction
$S=1+\sum_{i=1}^{\infty}t_{n}S^{n}.$ (1)
Let $S_{1}=\sum_{n=1}^{\infty}t_{n}$. Wildberger and Rubine [11, Theorem 12] showed that $S-1$ is divisible by $S_{1}$; i.e., there is a formal power series $G$ in the $t_{i}$, which they called the “Geode,” satisfying $S=1+GS_{1}$. (See also Rubine [8].)
@@
In Section 3 we give combinatorial interpretations to the series $S$, $G$, and $H$, and to the formulas relating them, in terms of lattice paths. We represent the paths by sequences of integers greater than or equal to $-1$, where the weight of $n\geq 0$ is $t_{n+1}$ and the weight of $-1$ is $1$. Then $S$ counts sequences with sum 0 and every partial sum nonnegative (excursions), $G$ counts sequences with every partial sum nonnegative, and $H$ counts sequences with every nonempty partial sum positive.
Wildberger and Rubine worked with polygon dissections rather than lattice paths. They did not include the variable $t_{1}$ as it is not so relevant to polygon dissections.
@@ 2. Formulas for the Geode
We first show the existence of a power series $G$ in the variables $t_{1},t_{2},t_{3},\dots$ satisfying $S=1+GS_{1}$.
@@
Then (4) follows directly from (6). Given (4), to prove (5) it is sufficient to show that $H^{-1}-G^{-1}=S_{1}$. From the definition of $H$, we have $H^{-1}=S/G$, so $H^{-1}-G^{-1}=(S-1)/G=S_{1}$. ∎
Equation (4) leads to a short proof of one of the conjectures of Wildberger and Rubine [11, Section 11]. Their conjecture is that
@@ 3.1. Paths and factorization
Next, we describe a factorization that will be useful in deriving the combinatorial interpretations for $S$, $G$ and $H$. The height of point $(m,n)$ is $n$, so the height of the endpoint of the path $(s_{1},\dots,s_{k})$ is $s_{1}+\cdots+s_{k}$. An excursion is a path (possibly empty) that ends at height 0 and never goes below height 0. (See Figure 2 for an example.)
@@
From here on, we assume that all steps are in the set $\{-1,0,1,2,\dots\}$. We call nonnegative steps up steps (so 00 is considered an up step) and we call $-1$ the down step. It will be convenient in describing factorizations of paths to denote the up step $n$ by $U_{n}$ and the down step $-1$ by $D$.
@@
Let $P$ be a path that ends at height $-n$, where $n\geq 0$, and never goes below height $-n$. Then $P$ can be factored uniquely as $E_{1}DE_{2}D\cdots DE_{n+1}$ where each $E_{i}$ is an excursion and $D$ is the down step $-1$.
@@
We proceed by induction on $n$. If $n=0$ then $P$ is an excursion. If $n>0$ then $P$ must have a first down step from height 0 to height $-1$. (This is where we use the assumption that there are no steps less than $-1$.) Cutting $P$ before and after this down step shows that $P=E_{1}DQ$ where $Q$, when translated to start at the origin, ends at height $-(n-1)$ and never goes below height $-(n-1)$. We then apply induction to $Q$. We leave it to the reader to verify that this factorization is unique. ∎
In Lemma 3.1, and similarly in other factorization statements, it is understood that the converse is also true, in the sense that if $P$ is a path with a factorization of the indicated form then $P$ is a path that ends at height $-n$ and never goes below height $-n$.
@@
Let $P$ be a path that ends at height $-n$, where $n\geq 0$, and stays above height $-n$ until the end. If $n>0$ then $P$ can be factored uniquely as $E_{1}DE_{2}D\cdots E_{n}D$ where each $E_{i}$ is an excursion and $D$ is the down step $-1$, and if $n=0$ then $P$ is the empty path.
@@
The case $n=0$ is evident, so suppose that $n>0$. Since $P$ satisfies the hypothesis of Lemma 3.1, it can be factored uniquely as $E_{1}DE_{2}D\cdots DE_{n+1}$ as in Lemma 3.1. But since $P$ stays above height $-n$ until the end, it must end with a down step, so $E_{n+1}$ is empty. ∎
@@ 3.2. Generating functions
We call a path nonnegative if it never goes below the $x$-axis; i.e., if every partial sum is nonnegative. (Nonnegative paths are sometimes called meanders.) We call a path positive if it stays above the $x$ axis; i.e., every nonempty partial sum is positive. Note that the empty path is a positive path.
@@
The power series $H$ is the generating function for positive paths.
@@ 3.3. Free monoids
Free monoids are useful in working with generating functions for the following reason. Suppose that $M$ is a free monoid. Suppose that a weight function is defined on $M$ so that the weight of an element of $M$ is the product of the weights of its constituent primes. Then if $U$ is the sum of the weights of all elements of $M$ and $V$ is the sum of the weights of all the primes of $M$, then
@@
Then we have the following lemma due to Schützenberger [9, Theorem 1.4]; see also Gessel and Li [4, Lemma 3] for a conveniently accessible proof. (We will use only the sufficiency of Schützenberger’s criterion.)
@@
We are now ready to explain the combinatorial interpretations of the formulas of Theorem 2.1.
For (2) we need to show that the generating function for arches is $\sum_{n=0}^{\infty}t_{n+1}S^{n}$. The proof is similar to the proof of Theorem 3.3. An arch must start with an up step $n$ for some $n\geq 0$. Then by Lemma 3.2, it can be factored as $U_{n}E_{1}DE_{2}D\cdots E_{n}D$. This gives the generating function for arches as $\sum_{n=0}^{\infty}t_{n+1}S^{n}$.
@@
For (4) we need to describe a decomposition for prime nonnegative paths. Let $P$ be a prime nonnegative path and suppose that $P$ starts with the up step $n$ so that $P$ may be factored as $U_{n}Q$ for some path $Q$. Then $P$ ends at height less than or equal to $n$ since if not, the factors $U_{n}$ and $Q$ would both be nonnegative and nonempty so $P$ would not be a prime. Suppose that $P$ ends at height $j$, where $0\leq j\leq n$. Then there cannot be any points of $P$ at a height less than or equal to $j$ other than the starting and ending points, since otherwise cutting $P$ at such a point would factor it into two nonempty nonnegative paths. Thus we may apply Lemma 3.2 to $Q$ to see that $Q$ can be factored uniquely as $E_{1}DE_{2}D\cdots E_{n-j}D$ where each $E_{i}$ is an excursion. (If $j=n$ then $Q$ is the empty path.) Thus the generating function for prime nonnegative paths with these values of $n$ and $j$ is $t_{n+1}S^{n-j}$. Summing over $j$ from 0 to $n$ gives $t_{n+1}(1+S+S^{2}+\cdots+S^{n})$, so the generating function for prime nonnegative paths is $\sum_{n=0}^{\infty}t^{n+1}(1+S+S^{2}+S^{n})$.
For (5) we describe a decomposition for prime positive paths. We omit the justification, as it is similar to that for prime nonnegative paths. Every prime positive path starting with the up step $n$ can be uniquely factored as $U_{n}E_{1}DE_{2}D\cdots DE_{n-j}$, where each $E_{i}$ is an excursion and $1\leq j\leq n$. Thus the generating function for prime positive paths is $\sum_{n=1}^{\infty}t_{n+1}(S+S^{2}+\cdots+S^{n})$.

### v2

@@ Abstract.
Let $t_{1},t_{2},\dots$ be variables, and let $S$ be the formal power series in the variables $t_{1},t_{2},\dots$ satisfying $S=1+\sum_{n=1}^{\infty}t_{n}S^{n}.$ Let $S_{1}=\sum_{n=1}^{\infty}t_{n}$. Wildberger and Rubine recently showed that there is a formal power series $G$ in the $t_{i}$, which they called the “Geode,” satisfying $S=1+GS_{1}$. In this paper we discuss some of the properties of the Geode and of the related series $H=G/S$, which satisfies $S=1/(1-HS_{1})$. We show that
@@ 1. Introduction
$S=1+\sum_{n=1}^{\infty}t_{n}S^{n}.$ (1)
Let $S_{1}=\sum_{n=1}^{\infty}t_{n}$. Wildberger and Rubine [14, Theorem 12] showed that $S-1$ is divisible by $S_{1}$; i.e., there is a formal power series $G$ in the $t_{i}$, which they called the “Geode,” satisfying $S=1+GS_{1}$. (See also Rubine [11].)
@@
In Section 3 we give combinatorial interpretations to the series $S$, $G$, and $H$, and to the formulas relating them, in terms of lattice paths. (Wildberger and Rubine worked with polygon dissections rather than lattice paths.) We represent the paths by sequences of integers greater than or equal to $-1$, where the weight of $n\geq 0$ is $t_{n+1}$ and the weight of $-1$ is $1$. Then $S$ counts sequences with sum 0 and every partial sum nonnegative (excursions), $G$ counts sequences with every partial sum nonnegative, and $H$ counts sequences with every nonempty partial sum positive.
In Section 4 we look at what happens when we set $t_{n}=0$ for $n>2$. Here (1) becomes a quadratic equation which can be solved explicitly, and we obtain generating functins for Catalan, Motzkin, Riordan, and Schröder numbers.
In Section 5 we gave an indirect alternative proof of the lattice path interpretation of the Geode, using one of the standard techniques of lattice path enumeration.
@@ 2. Formulas for the Geode
We first show the existence of a power series $G$ in the variables $t_{1},t_{2},t_{3},\dots$ satisfying $S=1+GS_{1}$. (Wildberger and Rubine did not include the variable $t_{1}$ as it is not so relevant to polygon dissections.)
@@
Then (4) follows directly from (6).. Given (4), to prove (5) it is sufficient to show that $H^{-1}-G^{-1}=S_{1}$. From the definition of $H$, we have $H^{-1}=S/G$, so $H^{-1}-G^{-1}=(S-1)/G=S_{1}$. ∎
Equation (4) leads to a short proof of one of the conjectures of Wildberger and Rubine [14, Section 11]. Their conjecture is that
@@ 3.1. Paths and factorization
The reversal of a path $(s_{1},\dots,s_{k})$ is the path $(-s_{k},\dots,-s_{1})$. Geometrically, reversing a path is reflecting it in a vertical axis. We call a path nonnegative if it never goes below the $x$-axis; i.e., every partial sum is nonnegative. (Nonnegative paths are sometimes called meanders.) We call a path positive if it stays above the $x$ axis; i.e., every nonempty partial sum is positive. Note that the empty path is a positive path. We call a path reverse-nonnegative if it is the reversal of a nonnegative path and reverse-positive if it is the reversal of a positive path. (See Figure 2.) Then reverse-nonnegative paths are paths whose endpoints are lowest points, and reverse-positive paths are paths in which the endpoint is strictly lower than every other point.
Figure 2. The nonnegative path $(1,1,-2,1)$ and its reversal $(-1,2,-1,-1)$.
The height of point $(m,n)$ is $n$, so the height of the endpoint of the path $(s_{1},\dots,s_{k})$ is $s_{1}+\cdots+s_{k}$. An excursion is a path (possibly empty) that ends at height 0 and never goes below height 0. (See Figure 3.)) Thus an excursion is a path that is both nonnegative and reverse-nonnegative.
@@
From here on, we assume that all steps are in the set $\{-1,0,1,2,\dots\}$. We call nonnegative steps up steps (so 00 is considered an up step) and we call $-1$ the down step. It will be convenient in representing factorizations of paths by formulas to denote the up step $n$ by $U_{n}$ and the down step $-1$ by $D$. The restriction on the step set implies that reverse-nonnegative and reverse-positive paths have very simple decompositions, but nonnegative and positive paths, which are closely related to the Geode, are more complicated.
Next, we describe the decompositions for reverse-nonnegative and reverse-positive paths that are helpful in deriving the combinatorial interpretations for $S$, $G$ and $H$.
@@
Let $P$ be a reverse-nonnegative path that ends at height $-n$. Then $P$ can be factored uniquely as $E_{1}DE_{2}D\cdots DE_{n+1}$ where each $E_{i}$ is an excursion and $D$ is the down step $-1$.
@@
We proceed by induction on $n$. If $n=0$ then $P$ is an excursion. If $n>0$ then $P$ must have a first down step from height 0 to height $-1$. (This is where we use the assumption that there are no steps less than $-1$.) Cutting $P$ before and after this down step shows that $P=E_{1}DQ$ where $Q$ is a reverse-nonnegative path that, when translated to start at the origin, ends at height $-(n-1)$. We then apply induction to $Q$. We leave it to the reader to verify that this factorization is unique. ∎
@@
In Lemma 3.1, and similarly in other factorization statements, it is understood that the converse is also true, in the sense that if $P=E_{1}DE_{2}D\cdots DE_{n+1}$ where each $E_{i}$ is an excursion, then $P$ is a reverse-nonnegative path that ends at height $-n$.
@@
Let $P$ be a reverse-positive path that ends at height $-n$. Then $P$ can be factored uniquely as $E_{1}DE_{2}D\cdots E_{n}D$ where each $E_{i}$ is an excursion and $D$ is the down step $-1$. (If $n=0$ this means that $P$ is the empty path.)
@@
The case $n=0$ is clear, so suppose that $n>0$. Since $P$ satisfies the hypothesis of Lemma 3.1, it can be factored uniquely as $E_{1}DE_{2}D\cdots DE_{n+1}$ as in Lemma 3.1. But since $P$ stays above height $-n$ until the end, it must end with a down step, so $E_{n+1}$ is empty. ∎
@@ 3.2. Generating functions
The series $H$ is the generating function for positive paths.
@@ 3.3. Free monoids
Free monoids are useful in working with generating functions for the following reason. Let $M$ be a free monoid. Suppose that a weight function is defined on $M$ so that the weight of an element of $M$ is the product of the weights of its constituent primes. Then if $U$ is the sum of the weights of all elements of $M$ and $V$ is the sum of the weights of all the primes of $M$, then
@@
Then we have the following lemma due to Schützenberger [12, Theorem 1.4]; see also Gessel and Li [7, Lemma 3] for a conveniently accessible proof. (We will use only the sufficiency of Schützenberger’s criterion.)
@@
Similarly, the set of reverse-nonnegative paths and the set of reverse-positive paths are free monoids. Lemma 3.2 shows that the prime reverse-positive paths are the paths of the form $ED$, where $D$ is an excursion. On the other hand, Lemma 3.1 does not describe the prime reverse-nonnegative paths; we leave it to the curious reader to determine what they are.
We are now ready to explain the combinatorial interpretations of the formulas of Theorem 2.1. For (2) we need to show that the generating function for arches is $\sum_{n=0}^{\infty}t_{n+1}S^{n}$. The proof is similar to the proof of Theorem 3.3. An arch must start with an up step $n$ for some $n\geq 0$, followed by a reverse-positive path ending at height $-n$. Then by Lemma 3.2, it can be factored as $U_{n}E_{1}DE_{2}D\cdots E_{n}D$. This gives the generating function for arches as $\sum_{n=0}^{\infty}t_{n+1}S^{n}$.
@@
For (4) we need to describe a decomposition for prime nonnegative paths. Let $P$ be a prime nonnegative path and suppose that $P$ starts with the up step $n$ so that $P$ may be factored as $U_{n}Q$ for some path $Q$. Then $P$ ends at height less than or equal to $n$ since if not, the factors $U_{n}$ and $Q$ would both be nonnegative and nonempty so $P$ would not be a prime. Suppose that $P$ ends at height $j$, where $0\leq j\leq n$. Then there cannot be any points of $P$ at a height less than or equal to $j$ other than the starting and ending points, since otherwise cutting $P$ at such a point would factor it into two nonempty nonnegative paths. Thus $Q$ is a reverse-positive path, and we may apply Lemma 3.2 to $Q$ to see that $Q$ can be factored uniquely as $E_{1}DE_{2}D\cdots E_{n-j}D$ where each $E_{i}$ is an excursion. (If $j=n$ then $Q$ is the empty path.) Thus the generating function for prime nonnegative paths with these values of $n$ and $j$ is $t_{n+1}S^{n-j}$. Summing over $j$ from 0 to $n$ gives $t_{n+1}(1+S+S^{2}+\cdots+S^{n})$, so the generating function for prime nonnegative paths is $\sum_{n=0}^{\infty}t_{n+1}(1+S+S^{2}+S^{n})$.
For (5) we describe a decomposition for prime positive paths. We omit details, as this decomposition is similar to that for prime nonnegative paths. Every prime positive path starting with the up step $n$ ends at height $j$ for some $j$ with $1\leq j\leq n$ and can be factored as $U_{n}Q$, where $Q$ is reverse-nonnegative path. By Lemma 3.1, $Q$ can be factored as $E_{1}DE_{2}D\cdots DE_{n-j+1}$, where each $E_{i}$ is an excursion. Thus the generating function for prime positive paths is $\sum_{n=1}^{\infty}t_{n+1}(S+S^{2}+\cdots+S^{n})$.
@@ 4. Catalan, Motzkin, Riordan, and Schröder numbers
5. Appendix: Another proof of Theorem 3.4
In this section we present another proof of Theorem 3.4, which states that the Geode $G$ is the generating function for nonnegative paths. This proof is less straightforward than the one in Section 3 but is an interesting application of lattice path enumeration techniques. It proceeds by first using an indirect method to count nonnegative paths with a nontrivial weight on down steps, and then carefully setting the weight for down steps to $1$.
We start by recalling a well-known factorization for lattice paths, sometimes called the Wiener-Hopf factorization, which applies to paths with arbitrary steps. We can factor a path into three parts (each possibly empty) by cutting the path at the first lowest point and at the last lowest point. The first part is a reverse-positive path, the second part is an excursion, and the third part is a positive path. An example is shown in Figure 8. This factorization is quite useful in lattice path enumeration. See [2, 3, 4, 6, 9] for some applications.
Figure 8. The Wiener-Hopf factorization of a path
For our problem, we want to combine the second and third parts into a nonnegative path, so we have a factorization of every path into a reverse-positive path followed by a nonnegative path.
The set of all paths is easy to count, and Lemma 3.2 allows us to count reverse-positive paths. However, there is a complication. There is no problem in assigning down steps the weight 1 when counting nonnegative paths, since for any finite multiset of up steps, there are only finite many nonnegative paths using these up steps, no matter how many down steps they have. But for arbitrary paths, and for reverse-positive paths, there are infinitely many paths with a given multiset of up steps. So to get a generating function identity from the factorization of arbitrary paths into reverse-positive paths and nonnegative paths, we change our definition of the generating function for a set of paths: we keep the weight of $t_{i+1}$ for the up step $i\geq 0$, but we change the weight of the down step $-1$ from 1 to a new variable $y$. Let $S(y)$ be the new generating function for excursions, so $S(1)=S$ by Theorem 3.3. Let $N(y)$ be the new generating function for nonnegative paths. Then to prove Theorem 3.4 we will show that $N(1)=G$.
By Lemma 3.2 the new generating function for reverse-positive paths is $\bigl{(}1-S(y)y\bigr{)}^{-1}$. The new generating function for all paths is $(1-y-t_{1}-t_{2}-\cdots)^{-1}=(1-y-S_{1})^{-1}$. Then the Wiener-Hopf factorization gives the identity
$\frac{1}{1-y-S_{1}}=\frac{1}{1-S(y)y}N(y).$ (12)
We would like to determine $N(1)$ from (12). We cannot set $y=1$ in (12), since the result would not converge as a formal power series in $t_{1},t_{2},\dots$. But this problem is easy to fix. We replace (12) with the equivalent equation
$1-S(y)y=(1-y-S_{1})N(y).$
Now there is no problem in setting $y=1$ and we find that $1-S=-S_{1}N(1)$, so $N(1)=(S-1)/S_{1}=G$, as desired.
Acknowledgment. I would like to thank Dean Rubine and Martin Rubey for pointing out typos in earlier versions of this paper.